OpenStudy (anonymous):

Solve the given equation. If necessary, round to four decimal places. 8^2x = 11

OpenStudy (anonymous):

take the log of both sides

OpenStudy (anonymous):

log(8^2x)=log(11)

OpenStudy (anonymous):

rule with logs is that an exponent can be moved to the front so: (2x)*log(8)=log(11)

OpenStudy (anonymous):

2x=(log(11))/(log(8))

OpenStudy (anonymous):

Ok Thank You!! it makes more sense now!

OpenStudy (anonymous):

no problem